OpenStudy (anonymous):

Solve the following quadratic equations. (hint: use factorization) HELP ME QUICK! please n^2+3n-12=6

OpenStudy (aroub):

n^2+3n-12-6=0 n^2+3n-18=0 (n+6)(n-3)=0 n=-6 , n=3

OpenStudy (anonymous):

Thanks Thanks... would you help me with another one?

OpenStudy (aroub):

Sure!

OpenStudy (anonymous):

THANKS!!! -4k^2-8k-3=-3-5k^2

OpenStudy (aroub):

-4k^2+5k^2-8k-3+3=0 k^2-8k=0

OpenStudy (aroub):

I don't think you can factor it more..

OpenStudy (aroub):

or only : k(k-8)=0

OpenStudy (aroub):

So, k=0 , k=8

OpenStudy (anonymous):

Its alright.., would you actually help me with my homework for tomorrow? :( I have lot of questions

OpenStudy (aroub):

I'll try my best =)

OpenStudy (anonymous):

thanks :) so the next one is Solve for x: x^2+8x=0

OpenStudy (aroub):

x=-8 and x=0